OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (tkhunny):

Have you considered calculating the slope of the line connecting the two points?

OpenStudy (anonymous):

yes @tkhunny i got 1x+y=3 and it was wrong

OpenStudy (anonymous):

@thomaster

OpenStudy (tkhunny):

E(-2, 2), F(5, 1) Well, you made some error along the way. Keep trying until you don't make errors. m = (2-1)/(-2-5) = 1/(-7) = -1/7 y-2 = (-1/7)(x+2) 7y - 14 = -x - 2 x + 7y - 12 = 0 x + 7y = 12 One of those.